OpenStudy (anonymous):

at a concert,$720 was collected for hot dogs, hamburgers and soft drinks.all three items sold for $1 each.twice as many hot dogs were sold as hamburgers.three times as many soft drinks were sold as hamburgers.what was the number of soft drinks sold? help pls

OpenStudy (anonymous):

@lucaz

OpenStudy (lucaz):

hamburger=h, soft drink=s, hot=d $1(h)+$1(s)+$1(d)=$720 d=2h, and s=3h, subsitute in the equation h+3h+2h=720 6h=720 h=720/6=120 go back to s=3h and replace h by 120 to find number of soft drinks s=3(120)=360

OpenStudy (anonymous):

thank uu